# Twisted Superfields

### v8

@@ Appendix A Quantum Effective potential
Since they are in the $(3,\tilde{3})$ and $(\tilde{3},3)$ representations, the gauginos and the fermions from the twisted adjoint superfield contribute zero to both $T^{0}_{V_{R}+}$ and $T^{0}_{A_{R}+}$. In addition, for the second term of (A.4) they generate a factor of $f^{bad}d^{cad}$ which vanishes. Therefore, those fermions do not contribute to the axial anomaly.
@@
The anomaly-generating axial U(1) charges of the superfields $Q_{mF}$, $\tilde{Q}_{mF}$, and $\Phi$ are 1, 1, and 0, respectively. Therefore, the following factors are invariant to the U(3)$\times$×U(3) local gauge symmetry and obey the axial selection rule:
@@
where $n$ is a non-negative integer. In the second “meson” configuration above, the $m$ and $F$ indices are restricted to combinations that have the same charge for $V_{+}^{0}$ in eq (1.17). For example, $M_{2213}^{\left(n\right)}=0$ (since $q_{21}\neq q_{23}$), but $M_{2113}^{\left(n\right)}\neq 0$ (since $q_{21}=q_{13}$).
@@
$Q_{F}=Q_{1F}+Q_{2F}\hskip 28.45274pt\tilde{Q}_{F}=\tilde{Q}_{1F}+\tilde{Q}_{2F}$ (A.7)
$M_{FF^{\prime}}^{(n)}=\Lambda^{-1}\tilde{Q}_{F}\Phi^{n}Q_{F^{\prime}}\,.$
In the meson factors, any terms that are not gauge invariant are set to zero. For example,
$M_{12}^{(0)}=\Lambda^{-1}(\tilde{Q}_{11}Q_{12}+\tilde{Q}_{21}Q_{12})\,.$ (A.8)
The terms $\tilde{Q}_{11}Q_{22}=\tilde{Q}_{21}Q_{22}=0$ inside of $M_{12}^{(0)}$ since $q_{21}=q_{11}\neq q_{22}$.
The following determinant is then defined
$\Delta^{(pqr)}={\rm det}\left|\begin{array}[]{ccc}{M_{11}^{(p)}}&{M_{12}^{(q)}}&{M_{13}^{(r)}}\\ {M_{21}^{(p)}}&{M_{22}^{(q)}}&{M_{23}^{(r)}}\\ {M_{31}^{(p)}}&{M_{32}^{(q)}}&{M_{33}^{(r)}}\end{array}\right|\,$ (A.9)
Following symmetry arguments similar to those used for SQCD, quantum interactions will generate low energy effective superpotential terms with mass dimension 3 and positive powers of $\Lambda$ that are constructed from the above determinants as well as factors of ${\rm Tr}\left(\Phi^{r}\right)$, and other gauge invariant factors such as those mentioned later in this Appendix.
@@
$O_{0}=\left({\rm Tr}\left(\Phi^{2}\right)\right)^{2}\left(\Delta^{(000)}\right)^{-1/3}\,.$ (A.10)
The scalar potential can be derived from the superpotential terms like the one above by taking derivatives to extract “f terms” proportional to $\theta^{2}f$ and inserting them into the f-term part of eq (2.1). The quantum vacuum can be found by minimizing the vev of the scalar potential after including these quantum contributions.
@@
$\bar{O}_{0}=\left\langle O_{0}\right\rangle=\frac{\Lambda\left({\rm tr}\left\langle\varphi_{1}^{2}\right\rangle+{\rm tr}\left\langle\varphi_{2}^{2}\right\rangle\right)^{2}}{\left(\bar{\phi}_{21}\bar{\tilde{\phi}}_{21}\bar{\phi}_{22}\bar{\tilde{\phi}}_{22}\bar{\phi}_{23}\bar{\tilde{\phi}}_{23}\right)^{1/3}}.$ (A.11)
In section 2, it was noted that if the model includes large tree-level masses $\tilde{m}_{21}$ and $\tilde{m}_{22}$, then the vevs $\bar{\tilde{\phi}}_{21}$ and $\bar{\tilde{\phi}}_{22}$ vanish clasically. However, in order to stabilize expressions such as the one above, it is assumed that $\bar{\tilde{\phi}}_{21}$ and $\bar{\tilde{\phi}}_{22}$ acquire small vevs quantum mechanically. This is mentioned in section 3. As shown below, those quantum-generated small vevs lead to a very large right-chiral neutrino Majorana mass term that allows the model to reproduce observed neutrino masses and mixing.
@@
Minimizing the quantum scalar potential generates the adjoint vevs in $\left\langle\varphi_{2}\right\rangle$ of eq (3.4) to reduce ${\rm tr}\left\langle\varphi_{2}^{2}\right\rangle=-{\rm tr}\left\langle\varphi_{1}^{2}\right\rangle$ in scalar potential contributions proportional to $\bar{O}_{0}$. Since ${\rm tr}\left\langle\varphi_{1}^{2}\right\rangle\neq 0$ (see the beginning of section 3), minimization makes ${\rm tr}\left\langle\varphi_{2}^{2}\right\rangle\neq 0$ for the vacuum solution.
@@
Once the scalar vevs have been adjusted and a minimum of the quantum potential has been achieved, quantum superpotential terms can also produce fermion and scalar mass terms. For example, the term $O_{0}$ generates the following terms that modify the tree-level masses $m_{m11}$ and $m_{m22}$:
$-\tfrac{1}{3}\bar{O}_{0}\sum_{m}\left(\frac{\tilde{Q}_{m1}Q_{m1}}{\bar{\tilde{\phi}}_{21}\bar{\phi}_{21}}+\frac{\tilde{Q}_{m2}Q_{m2}}{\bar{\tilde{\phi}}_{22}\bar{\phi}_{22}}\right),$ (A.12)
where terms involving $1/\bar{\phi}_{23}$ and $1/\bar{\tilde{\phi}}_{23}$ have been ignored due to their inverse unification-scale vevs.
The quantum potential also generates a Majorana mass term along with its scalar mass counterpart:
$\frac{4\theta^{2}\bar{O}_{0}}{9\bar{\tilde{\phi}}_{21}^{2}}\left(2(\tilde{f}_{21})_{1}(\tilde{\phi}_{21})_{1}-\tilde{\nu}_{W}^{(1)}\tilde{\nu}_{W}^{(1)}\right)\,,$ (A.13)
@@
where terms involving $1/\bar{\phi}_{21}$ have been ignored. Since the vev $\bar{\tilde{\phi}}_{21}$ is very small, the second term generates a very large neutrino Majorana mass. The first term generates the corresponding scalar mass, where the notation $(\tilde{\phi}_{21})_{1}$ refers only to the first SU(3) component of that scalar.
The quantum-generated mass terms in eqs (A.12) and (A.13) were constructed in the context of the $m=1$ flavor $F=1$ being identified with the $m=2$ flavor $F=1$. Due to the $3\times 3$ flavor symmetry in the $m=1$ sector, there are also analogous terms where the $m=1$ flavors $F=2$ or $F=3$ are identified with the $m=2$ flavor $F=1$. With those included, the neutrino Majorana mass term becomes
$\frac{4\theta^{2}\bar{O}_{0}}{9\bar{\tilde{\phi}}_{21}^{2}}\sum_{F}\tilde{\nu}_{W}^{(F)}\tilde{\nu}_{W}^{(F)}\,.$ (A.14)
In other words, all three right-handed neutrinos have the same very large Majorana mass, $\tilde{M}_{\nu}$. For $m=2$ scalars, there is no analogous flavor rotation since the $m=2$ sector only has $1\oplus 2\times 2$ flavor symmetry. So only $(\tilde{\phi}_{21})_{1}$ gets a very large mass by the above mechanism.
Under the above $m=1$ flavor rotation, the $m=1$ mass terms of eq (A.12) become
$-\tfrac{1}{3}\bar{O}_{0}\left(\frac{1}{\bar{\tilde{\phi}}_{21}\bar{\phi}_{21}}+\frac{1}{\bar{\tilde{\phi}}_{22}\bar{\phi}_{22}}\right)\sum_{F}\tilde{Q}_{1F}Q_{1F}\,.$ (A.15)
The same mechanism that generates large Majorana neutrino masses can also generate an effective coupling important for scalar-mediated quark decays. Just as quantum effects generate $O_{0}$ in eq (A.10), they will also generate an $O_{2}$ that replaces $\Delta^{(000)}$ in $O_{0}$ with $\Delta^{(200)}+\Delta^{(020)}+\Delta^{(002)}$, while changing $({\rm Tr}(\Phi^{2}))^{2}\to({\rm Tr}(\Phi^{2}))^{7/3}$. This generates the following quantum superpotential term:
$-i\frac{\sqrt{2}\theta^{2}\bar{O}_{c}}{3c\bar{\varphi}_{2}^{2}\bar{\tilde{\phi}}_{21}}\left(\tilde{u}^{(1)}\varphi_{1}u_{W}^{A}\right)\,.$ (A.16)
@@
Just as for the Majorana neutrino masses, since $\bar{\tilde{\phi}}_{21}$ is very small, this quantum-generated coupling is large. This coupling is used in section 5 when showing how the model can reproduce $Z$ pole observables.
@@
The quantum superpotential can also include expressions like $O_{0}$ of eq (A.10) where the “meson” factors in the denominator are generalized in the following way:
$M^{\left(0\right)}_{m_{1}m_{2}FF^{\prime}}\to M^{\left(n\right)}_{m_{1}m_{2}FF^{\prime}}.$ (A.18)
These terms must have corresponding additional factors of $\Phi^{2}$ in the numerator. For example, a term where the flavor 2 and 3 mesons have a total of 3 additional factors of $\Phi^{2}$ in the denominator, there must be an additional factor of $\Phi^{2}$ in the numerator. One of these terms can have ${\rm Tr}\left(W^{0\alpha}W^{\prime}_{\alpha}\Phi^{3}\right)$ in the numerator. This term generates an adjoint-gaugino mass term for down-type quarks but not for up-type quarks (since the “11” component of $\langle\Phi^{2}\rangle$ is zero). The presence of terms like these provide justification for not forcing these types of terms to be the same in the up-type and down-type quark matrices of section 3. Terms involving ${\rm Tr}\left(\Phi^{4}\right)$ in the numerator can similarly lead to differences in adjoint-adjoint quark mass terms for up-type and down-type quarks.
Superpotential terms with ${\rm Tr}(\Phi^{2})^{3}$ in the numerator can lead to mass terms like that of eq (A.12), but where $\tilde{Q}_{m1}Q_{m1}\to\tilde{Q}_{m1}\langle\Phi^{2}\rangle Q_{m1}$. These terms only generate masses for the second two components of lepton triplets. This shows that the first component may have a different mass than the second two components. This freedom is used in the lepton mass matrices of section 3.
For superpotential terms in which the numerator has dimension 5 (such as $W^{0\alpha}W^{0}_{\alpha}{\rm Tr}\left(\Phi^{2}\right)$), there must be an odd number of $\Phi$ factors in the denominator. A term like this involving $M^{(1)}_{2111}$ generates a quark mass term for $\tilde{u}_{W}^{A}u^{(1)}$ that supplements the tree level term involving $\bar{\tilde{\phi}}_{21}\Gamma_{211}$. Similarly, many other quantum-generated couplings can supplement equivalent tree-level couplings.

### v9

@@ Appendix A Quantum Effective potential
Since they are in the (3,3~)($\tilde{3}$) and $(\tilde{3},3)$ representations, the gauginos and the fermions from the twisted adjoint superfield contribute zero to both $T^{0}_{V_{R}+}$ and $T^{0}_{A_{R}+}$. In addition, for the second term of (A.4) they generate a factor of $f^{bad}d^{cad}$ which vanishes. Therefore, those fermions do not contribute to the axial anomaly.
@@
[…]
@@
where $n$ is a non-negative integer. From eq (1.17), all of the “meson” configurations ($M$) above are invariant to nonAbelian supergauge transformations, but some are not invariant to Abelian supergauge transformations dependent on $\Lambda_{+}^{0}$. This is addressed below.
@@
$Q^{\prime}_{1F}=Q_{1F}+Q_{2F}\hskip 28.45274pt\tilde{Q}^{\prime}_{1F}=\tilde{Q}_{1F}+\tilde{Q}_{2F}$ (A.7)
$M_{1FF^{\prime}}^{(n)}=\Lambda^{-1}\tilde{Q}^{\prime}_{1F}\Phi^{n}Q^{\prime}_{1F^{\prime}}\,.$
The subscript “1” on $^{\prime}, Q~\tilde{Q}^{\prime}$, and $M$ denotes that the $m=1$ flavor $F=1$ was identified with the $m=2$ flavor $F=1$ in the above definitions. Next, as mentioned above, the $m=1$ flavors are redefined with $1\to 2\to 3\to 1$ (e.g. $Q^{\prime}_{21}=Q_{12}+Q_{21}$ or $Q^{\prime}_{31}=Q_{13}+Q_{21}$).
Just as flavor determinants are used in the quantum vacuum of SQCD, the following flavor determinant is defined:
$\Delta^{(pqr)}_{i}={\rm det}\left|\begin{array}[]{ccc}{M_{i11}^{(p)}}&{M_{i12}^{(q)}}&{M_{i13}^{(r)}}\\ {M_{i21}^{(p)}}&{M_{i22}^{(q)}}&{M_{i23}^{(r)}}\\ {M_{i31}^{(p)}}&{M_{i32}^{(q)}}&{M_{i33}^{(r)}}\end{array}\right|\,.$ (A.8)
As mentioned above, some of the $M$ factors are not invariant to Abelian supergauge transformations dependent on $\Lambda_{+}^{0}$.
Any terms in the determinant that are not completely supergauge invariant are set to zero. That means for example that in the term $M_{111}^{(0)}M_{123}^{(0)}M_{132}^{(0)}$,
$M_{111}^{(0)}\tilde{Q}_{12}Q_{23}\tilde{Q}_{23}Q_{22}=0\,$, (A.9)
since under a supergauge transformation, $\tilde{Q}_{12}Q_{23}\to\tilde{Q}_{12}Q_{23}e^{i\Lambda_{+}^{0}/\sqrt{12}}$ and there is nothing else in that term that can cancel the $\Lambda_{+}^{0}$ gauge dependence. On the other hand,
$M_{111}^{(0)}\tilde{Q}_{12}Q_{23}\tilde{Q}_{23}Q_{12}\neq 0\,$, (A.10)
since the $\tilde{Q}_{11}Q_{22}$ and $\tilde{Q}_{22}Q_{11}$ factors have $\Lambda_{+}^{0}$ gauge dependencies that cancel each other.
Following symmetry arguments similar to those used for SQCD, quantum interactions will generate low energy effective superpotential terms with mass dimension 3 and one power of $\Lambda$ in the numerator. These terms are constructed from the above flavor determinants as well as factors of $\left({\rm Tr}(\Phi^{2})\right)^{s}$ and other gauge invariant factors such as those mentioned later in this Appendix.
@@
$O_{0}=\left({\rm Tr}\left(\Phi^{2}\right)\right)^{2}\left(\tfrac{1}{3}\sum_{i}\Delta^{(000)}_{i}\right)^{-1/3}\,.$ (A.11)
The scalar potential can be derived from superpotential terms like the one above by taking Taylor-expansion derivatives to extract “f terms” proportional to $\theta^{2}f$ and inserting them into the f-term part of eq (2.1). The quantum vacuum can be found by minimizing the vev of the scalar potential after including these quantum contributions.
@@
$\bar{O}_{0}$ $=\left\langle O_{0}\right\rangle=\frac{\Lambda\bar{\varphi}_{1}^{4}}{\left(\bar{\phi}_{21}\bar{\tilde{\phi}}_{21}\bar{\phi}_{22}\bar{\tilde{\phi}}_{22}\bar{\phi}_{23}\bar{\tilde{\phi}}_{23}\right)^{1/3}}$ (A.12)
$\bar{\varphi}_{1}^{2}$ $={\rm tr}\left\langle\varphi_{1}^{2}\right\rangle\,$,
where it was argued at the beginning of section 3 that ${\rm tr}\left\langle\varphi_{1}^{2}\right\rangle\neq 0$ in the quantum vacuum. In section 2, it was noted that if the model includes large tree-level masses $m_{22}$, $\tilde{m}_{21}$ and $\tilde{m}_{22}$, then the vevs $\bar{\phi}_{22}$, $\bar{\tilde{\phi}}_{21}$ and $\bar{\tilde{\phi}}_{22}$ vanish classically. However, in order to stabilize expressions such as the one above, it is assumed that these vevs acquire small values quantum mechanically. This is mentioned in section 3. As shown below, those quantum-generated small vevs lead to very large right-chiral neutrino Majorana mass terms that allow the model to reproduce observed neutrino masses and mixing.
@@
Once the scalar vevs have been adjusted and a minimum of the quantum potential has been achieved, second-derivative Taylor-expansion terms of the quantum superpotential can also produce fermion and scalar mass terms.
For example, taking derivatives with respect to $\tilde{Q}^{\prime}_{i1}$ and $\tilde{Q}^{\prime}_{j1}$, evaluating that expression at its vev, then multiplying by $\tilde{Q}^{\prime}_{i1}$ and $\tilde{Q}^{\prime}_{j1}$ generates the following mass terms
$\frac{4\theta^{2}\bar{O}_{0}}{9\bar{\tilde{\phi}}_{21}^{2}}\left(2(\tilde{f}_{21})_{1}(\tilde{\phi}_{21})_{1}-\tfrac{1}{9}\sum_{FF^{\prime}}\tilde{\nu}_{W}^{(F)}\tilde{\nu}_{W}^{(F^{\prime})}\right)\,$, (A.13)
@@
where $2\theta\tilde{\nu}_{W}^{(F)}\theta\tilde{\nu}_{W}^{(F^{\prime})}=-\theta^{2}\tilde{\nu}_{W}^{(F)}\tilde{\nu}_{W}^{(F^{\prime})}$ was used. Since the quantum-generated vev $\bar{\tilde{\phi}}_{21}$ is very small, the second term generates a very large neutrino Majorana mass matrix. The first term generates the corresponding scalar mass, where the notation $(\tilde{\phi}_{21})_{1}$ denotes the first SU(3) component of the $\tilde{\phi}_{21}$ scalar triplet.
Taking derivatives of $O_{0}$ with respect to $\tilde{Q}^{\prime}_{i1}$ and $Q^{\prime}_{j1}$ generates the following mass terms:
$-\frac{\bar{O}_{0}}{3\bar{\tilde{\phi}}_{21}\bar{\phi}_{21}}\left(\tilde{Q}_{21}Q_{21}-\tfrac{4}{3}(\tilde{Q}_{21})_{4}(Q_{21})_{4}\right)$ (A.14)
$+\frac{\bar{O}_{0}\theta^{2}}{9\bar{\tilde{\phi}}_{21}\bar{\phi}_{21}}\sum_{F}\left(\tilde{\nu}_{W}^{(F)}\nu_{W}^{(F)}-\tfrac{4}{9}\sum_{F^{\prime}}\tilde{\nu}_{W}^{(F)}\nu_{W}^{(F^{\prime})}\right)\,$,
where a variant of eq (A.10) was used. The second line are neutrino Dirac masses that can form seesaws with the heavy neutrinos. This is discussed in the Neutrino part of section 3.
Taking derivatives with respect to $\tilde{Q}^{\prime}_{i2}$ and $Q^{\prime}_{j2}$ generates more mass terms:
$-\frac{\bar{O}_{0}}{3\bar{\tilde{\phi}}_{22}\bar{\phi}_{22}}\left(\tilde{Q}_{22}Q_{22}-\tfrac{4}{3}(\tilde{Q}_{22})_{5}(Q_{22})_{5}\right)$ (A.15)
$+\frac{\bar{O}_{0}\theta^{2}}{9\bar{\tilde{\phi}}_{22}\bar{\phi}_{22}}\sum_{F}\tilde{e}_{W}^{(F)}e_{W}^{(F)}\,.$
The masses in the second line form part of a seesaw to generate the observed charged lepton masses. Due to the applicable variant of eq (A.9), this term is diagonal in the charged lepton flavor (unlike neutrino mass terms).
Derivatives can also be taken with respect to adjoint fields. Taking derivatives with respect to the adjoint fermions $\chi$ and $\tilde{\chi}$ generates the following quark mass terms:
[…]
where the notation described below eq (3.16) for quark gauge eigenstates is being used. From the CKM fit to quark mass matrices in eq (3.16), $f^{\prime}_{L}=.99f_{L}+.14b_{L}+$.... So the above quantum generated term has a small $b\bar{b}$ component.
For each of the three cases above, if a third derivative is taken with respect to $Q_{21}$ before setting the expression to its vev, then $h_{t}$ Higgs interaction terms are generated. These are proportional to the expressions above but multiplied by $(\phi_{21})_{1}/\bar{\phi}_{21}\simeq\hat{g}(\phi_{21})_{1}/m_{t}$, where eq (3.12) was used. These terms allow $h_{t}$ to decay to leptons, to $c\bar{c}$ and to a lesser extent, to $b\bar{b}$. These decay mechanisms are mentioned in the Higgs decay part of section 5.
If a third derivative with respect to $(\varphi_{2})_{32}$ is taken instead, then new scalar interaction terms are generated for the scalar $(\varphi_{2})_{32}$. These terms are proportional to the expressions above but multiplied by $2(\varphi_{2})_{32}\bar{\varphi}_{2}/\bar{\varphi}_{1}^{2}$. So they enable decays of $(\varphi_{2})_{32}$ to leptons, $c\bar{c}$ and to a lesser extent, to $b\bar{b}$.
The operator $O_{0}$ could be generalized by inserting various powers of adjoint fields in the flavor determinant. One simple example generalization is
$\Delta^{(000)}_{i}\to\Delta^{(000)}_{i}+\sum_{n}\alpha_{n}\Delta^{(00n)}_{i}/\left({\rm Tr}\left(\Phi^{2}\right)\right)^{n/2}\,$, (A.17)
where $\alpha_{n}$ are constants. Of course this could be generalized further by placing various multiples of adjoint fields in different parts of the determinant. These additional terms not change the vev $\bar{O}_{0}$, but they do generate additional mass terms.
For example, with $n=1$, if derivatives are taken with respect to $\tilde{Q}_{21}$ and $(\varphi_{2})_{k1}$ then set to its vev, then multiplied by those fields, the following scalar mass terms are generated:
$\frac{\bar{O}_{0}\theta^{2}}{3\bar{\varphi}_{1}\bar{\tilde{\phi}}_{21}}\left((\tilde{f}_{21})_{k}(\varphi_{2})_{k1}+(\tilde{\phi}_{21})_{k}(f_{2})_{k1}\right)$ (A.18)
Solving the equations of motion for $(\tilde{f}_{21})_{2}$ and $(\tilde{f}_{21})_{3}$, this term generates large masses for $(\varphi_{2})_{21}$ and $(\varphi_{2})_{31}$. It does not cause a large mass for $(\varphi_{2})_{11}$, since the equation of motion for $(\tilde{f}_{21})_{1}$ is dominated by the much larger mass term of eq (A.13). The equations of motion for $(f_{2})_{21}$ and $(f_{2})_{21}$ generate those same large masses for $(\tilde{\phi}_{21})_{2}$ and $(\tilde{\phi}_{21})_{3}$.
Since these masses are generated at a scale much smaller than the unification scale and they only involve the $m=2$ sector, they could be generated by an effective theory after the $m=2$ SU(3) has been broken down to SU(2). In that case, the mass terms for $(\varphi_{2})_{21}$ and $(\tilde{\phi}_{21})_{2}$ do not need to be the same as those for $(\varphi_{2})_{31}$ and $(\tilde{\phi}_{21})_{3}$. In section 5, it is estimated that the mass of $(\tilde{\phi}_{21})_{2}$ is $\sim$3.5 TeV and that of $(\varphi_{2})_{31}$ is larger, >8>8 TeV.
There is another consequence of the mass generation happening in the sub-unification SU(2) theory. Since SU(2) is a self-conjugate group, a transformation generates the additional mass term $(\varphi_{2})_{13}(f_{21})_{3}+(f_{2})_{13}(\phi_{21})_{3}$. This results in the same large mass (>8>8 TeV) for the scalars $(\varphi_{2})_{13}(f_{21})_{3}$ and $(\phi_{21})_{3}$.
By taking $n=2$ in eq (A.17), the following quantum superpotential term is generated:
$i\frac{\sqrt{2}\theta^{2}\bar{O}_{c}}{3\bar{\varphi}_{1}^{2}\bar{\tilde{\phi}}_{21}}\left(\tilde{u}^{(1)}\varphi_{1}u_{W}^{A}\right)\,.$ (A.19)
@@
Just as for the Majorana neutrino masses, since $\bar{\tilde{\phi}}_{21}$ is very small, this quantum-generated coupling is large. This coupling is used in section 5 when describing a way that the model could potentially reproduce $Z$ pole observables.